OpenStudy (anonymous):

Help ? write 4/5 x 3/8 in lowest terms please explain how you did it :D

OpenStudy (anonymous):

Is this the same? \[\frac{ 4 }{ 5 }*\frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

When multiplying fractions together, you multiplyt he numerators to each other, then multiply the denominators together so, \[\frac{ 4*3 }{ 5*8 }\]

OpenStudy (anonymous):

\[\frac{ 4 * 3 }{ 5 * 8 } = \frac{ 12 }{ 40 } = \frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

4/5 x 3/8 Since 4 and 8 can cancel each other, So 4 becomes 1. And 8 becomes 2. Therefore, It will be, 1/5 x 3/2 So the answer must be 3/10.

OpenStudy (anonymous):

Thank you all!